Determine the area of the region bounded by the given set of curves. x=e^y, x=1, y=1 and y=2


1
Expert's answer
2021-05-12T16:11:23-0400

"x=e^y,x=1,y=1,y=2"


Area under curve is "A=\int_a^bf(x) dx"


"y=1,y=2,x=1"


"F(x)\Rightarrow x=e^y\Rightarrow y=ln(x)"


Area "A=\int_1^2 ln(x)dx"

     

   "=[xln|x|-x]_1^2


\\
 =[2ln|2|-2-(1ln|1|-1)]


\\
 =2ln|2|-2+1
\\
 =2\times 0.693-1=0.386"